The line through the points (4,5) and (10, -2)
Answer by checkley75(3666)

SLOPE = (Y2-Y1)/(X2-X1)
(-2-5)/(10-4)
-7/6 IS THE SLOPE (m)
NOW USE OPNE SET OF THE X & Y VALUES IN THE LINE EQUATION Y=mX+b AND SOLVE FOR b
5=-7/6*4+b
5=-28/6+b
b=5+28/6
b=(30+28)/6
b=58/6
b=29/3 THUS THE LINE EQUATION IS
Y=-7/6X+29/3